``Solve using the multiplication principle. DOnt forget to preform a check.-7/8x=-9/10The solution is x=______Type an integer or fraction. Can you help me please. i have come up with 3 diffretn answers''.
I have come up with 2 answers can you help me please!!!

Answer by rfer(16322)

-7/8x=-9/10
x=(-9/10)/(-7/8)
invert second fraction and multiply
x=-9/10*-8/7
x=72/70
x=1 1/35